Fermi National Laboratory
 

Accelerator Update for July 17, 2006 - July 19, 2006

 

Monday July 17

The day shift began with Operations monitoring store 4838 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the NuMI horn cooling work continuing, and with MiniBooNE taking beam.

At 8:46 AM, a Booster RF station (BRF8) tripped off for the first of four times in about 45 minutes. It reset each time.

Booster experts reported at 9:36 AM that they were bypassing BRF17 and this would hold off beam for a while. RF experts investigated a BRF8 problem during this downtime.

Beam resumed at 10:14 AM. Pbar resumed stacking.

Operations established beam to the Meson MTest experiment at 10:28 AM.

A Booster expert brought BRF17 back on line at 11:25 AM.

At 1:29 PM, a Linac RF station (KRF5) tripped off on a coil comparator fault. Operators and Linac experts responded. Beam resumed three minutes later.

At 1:51 PM, BRF7 & 8 tripped off. Both stations reset.

At 1:55 PM, BRF4 tripped off. It reset.

At 1:57 PM, Operations reported that water was coming out from under a power supply equipment rack at the Tevatron (TeV) sector C1 service building. An expert investigated and reported that repairs would have to be made soon, but they could wait until the store was terminated.

At 2:05 PM, Operations reported a line voltage alarm for the low energy Linac. It appeared that the lab’s line voltage was sagging. A Linac expert compensated for this.

At 2:12 PM, Operations reported that the Recycler stash was lost. Operators began preparing MI for an access to investigate and repair the problem.

Operations prepared Linac for an access at 2:54 PM.

At 3:20 PM, Operations reported that the LCW temperature at TeV sector A3 was creeping up. An expert investigated and found the LCW pump off. This will need to be turned on in between stores. The temperature was 102 degrees F and the power supplies trip at 115 degrees F.

MIRF technicians reported at 4:20 PM that they needed an hour to replace a tube inside the MIRF18 modulator.

Recycler and MI experts came out of access and Operations turned MI on at 5:35 PM.

Operations resumed sending beam to MTest at 6:29 PM.

Pbar resumed stacking at 6:33 PM.

Recycler and MI experts reported at 6:51 PM that the stash was lost due to a vacuum burst in the extraction line between MI and Recycler. The vacuum degraded and beam valves closed. Unfortunately, vacuum techs could not find a cause. The Run Coordinator decided that the Recycler would run for a couple of hours without antiprotons, but use reverse protons to conduct emittance measurements.

At 7:50 PM, LRF5 tripped off. After resetting many different pieces of equipment, operators discharged the crowbar circuit and then the reset worked. The station was down for thirty minutes.

MTest experimenters gave up for the night at 9:01 PM.

Operations prepared the Recycler around 9:10 PM to stash antiprotons. The stash ended and stacking resumed sometime around 10:19 PM.

At 10:46 PM, lightning caused the MI power supply loop interlock to trip. It reset.

Tuesday July 18

The midnight shift began with Operations monitoring store 4838 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 2:25 AM, Operations reported that the antiproton stack rate was bad. Operations called in an expert.

At 5 AM, Pbar stopped stacking to send antiprotons to the Recycler, but there were severe problems with the reverse proton tune up. Another expert came in to help. Pbar resumed stacking at 6:05 AM.

The Recycler began preparing to stash at 7:13 AM.

At 7:15 AM, LRF5 tripped off. Operators had the station back up within five minutes.

At 7:44 AM, a Booster vacuum pump tripped off and caused BRF7 & 8 to trip. Everything reset okay.

At 8:07 AM, BRF4 tripped off. It reset.

The stash ended and Pbar resumed stacking at 8:51 AM.

Operations established beam to MTest at 9:18 AM.

TeV experts began an end-of-store study at 11:03 AM.

The Recycler began preparing to stash at 12:55 PM. The stash ended around 2 PM.

The TeV end-of-store study ended and Operations terminated store 4838 at 2:24 PM. MI experts began a BPM study.

Operations put the TeV into shot setup at 3:02 PM.

MI completed their BPM study at 4:32 PM.

At 4:37 PM, Operations reported that the MI BPMs were not working and was slowing the reverse proton tune up. Experts investigated.

Operations established store 4839 at 6:23 PM with an initial luminosity of 102.29E30. Pbar resumed stacking.

MTest resumed getting beam at 6:18 PM.

Operations reported problems with QXR at 7 PM. At 7:42 PM, Operations called experts and then made adjustments. MTest reported that the beam was better at 8:07 PM.

The Recycler began preparing to stash at 9:19 PM. The stash ended around 10 PM.

At 10:58 PM, BRF17 tripped off for the first of three times in eighteen minutes. It reset okay.

Wednesday July 19

The midnight shift began with Operations monitoring store 4839 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 12:36 AM, BRF17 tripped off. Operators bypassed the station.

The Recycler began stashing antiprotons around 2 AM. Pbar resumed stacking at 2:54 AM.

Operations reported at 4:56 PM, that the Pbar LCW temperatures were rising. The Duty Mechanic investigated and said that everything was running nominally.

The Recycler began preparing to stash at 6:32 AM. The first transfer failed. The second transfer was better, 88% efficiency. Pbar resumed stacking at 6:51 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/19/2006    email Fermilab